# Decay of solutions of nonlinear Dirac equations

### v2

@@ Abstract.
We study the long-time behavior of small and large solutions to a broad class of nonlinear Dirac-type equations. Our results are classified in 1D massless and massive cases, 3D general and $n$ dimensional in generality. In the 1D massless case we prove that any globally defined solution converges to zero as time tends to infinity, within a spatial region expanding at a rate proportional to $t\log^{-2}t$. This result holds without assumptions on the smallness of initial data or specific power of nonlinearity, ruling out the existence of standing breather-like or solitary wave structures in this regime. In the 1D massive case, solitary waves are known to exist. Introducing new virial identities adapted to the Dirac’s distinctive algebra, we prove that there are “holomorphic” odd nonlinearities under which globally defined small odd solutions decay to zero on spatial compact sets as time tends to infinity. This result is extended to the 3D case under boundedness of the $H^{1}$ norm but without requiring the parity condition on the data, giving decay proofs for an important class of nonlinear Dirac models, and opening the door to the future use of virial identities to prove asymptotic stability of well-chosen Dirac solitary waves.
@@ 1.1. Setting
The (linear) Dirac equation was introduced as a relativistic version of the Schrödinger equation [32], being one of the most relevant models in relativistic quantum mechanics. It describes the self-interaction of Dirac fermions [32, 64]. Several authors have introduced and justified a nonlinear version of this model, see e.g. Soler and Thirring models [62, 65, 1]. Generally, the nonlinear Dirac model is
@@
Here $\psi:\mathbb{R}\times\mathbb{R}^{n}\to\mathbb{C}^{N}$ is a spinor-valued wave function, $m\in\mathbb{R}$ is the mass, and $N=\lfloor(2n+1)/2\rfloor$, where the summation convention is $\gamma^{\mu}\partial_{\mu}=\sum_{\mu=0}^{n}\gamma^{\mu}\partial_{\mu}$, and $\partial_{0}\equiv\partial_{t}$. The Dirac matrices $\gamma^{\mu}$ are chosen such that they satisfy the anti-commutativity property
@@
and where $\eta^{\mu\nu}$ is related to the Minkowski metric $\eta=\mbox{diag}(1,-1,\dots,-1)$. Some members of the Dirac family equation are invariant under Lorentz boosts, e.g. the Soler and Thirring models.
@@ 1.2. Well-posedness
The well-posedness problem of the nonlinear Dirac equation is challenging due to the lack of a defined sign on the Hamiltonian. Most works address the massive and the massless cases separately. The literature is extensive across both 1D and higher dimensions; here, we highlight most relevant works.
In the 1D case, Candy [18] used null coordinates to prove global existence in $H^{s}$ for all $s\geq 0$ in the case of the $L^{2}$-critical and integrable massive Thirring model. Pelinovsky [58] showed the existence of small-$H^{1}$-norm solutions for more general nonlinear Dirac equations with cubic and higher-order nonlinear terms, and also established scattering for global solutions with small initial data in $H^{1}\cap L^{1}$. In [46] the authors proved local well-posedness in $H^{s}$ with $s>-1/2$ and ill-posedness in $H^{-1/2}$ for the quadratic nonlinear Dirac equation. See [18, 58, 46, 54, 59, 7, 10, 60] and references therein for further details.
@@ 1.3. Main results
In this paper, our main goal is to study the long time asymptotic behavior of globally defined nonlinear Dirac solutions, and under minimal assumptions, and including integrable and nonintegrable models. Highly relevant for us will be the 1D and 3D cases. We first consider the 1D Dirac equation in laboratory coordinates:
@@
$U_{\omega}(x)=\frac{\gamma}{\sqrt{1+\omega}\cosh(\gamma x)+i\sqrt{1-\omega}\sinh(\gamma x)},\quad\gamma:=\sqrt{1-\omega^{2}}.$
These (standing) solitary waves are clear obstruction to decay, having interesting symmetries. Indeed, one can see that among the symmetries preserved by (1.6) one can find the (even + $i$ odd) and (odd + $i$ even) ones, which are precisely present in the solitary waves (1.7). Moreover, notice that the classical (odd + $i$ odd) symmetry, natural in scalar field models with odd nonlinearities, is not preserved this time. This makes the exercise of finding decay in the energy space in the 1D case highly a complicated task. Indeed, we shall concentrate efforts in another manifold of initial data not producing solutions such as (1.7). Within this region we shall obtain energy local decay.
@@
[…]
@@
It is well-known that the cubic Nonlinear Schrödinger (NLS) equation admits breather solutions. These solutions are characterized by being nontrivially spatially localized and periodic in time, distinguishing them from standing waves. Breather solutions possess a more intricate analytical structure; for example, they may involve complex interactions between different modes, which result in oscillatory behavior that is both spatially and temporally localized. One of the well-known breather for the focusing cubic NLS is the Satsuma-Yajima breather solution (see [4, 2] and the reference therein). For the Klein-Gordon equation (and for a class of 1D nonlinear wave equations), in [44], the authors proved the nonexistence of small odd breather solutions (see also [61]).
Similarly, Martínez [48] showed that, for a wide range of nonlinearities, any global odd solution to the 1D NLS equation decays to zero in compact regions of space as time tends to infinity, assuming only initial data in the energy space. Previous work addressed the decay/scattering problem under the condition that the power $p$ satisfies $53<p<$ 5, but this result improved the range including $1<p\leq 3$. More recently, [2] characterized the nonexistence of breather solutions for the $n$-dimensional NLS using the energy and power of the corresponding nonlinearity.
Given the similarities between the Dirac equation and both the Klein-Gordon and the classical Nonlinear Schrödinger equations, a natural question arises: Do breather solutions exist for the Dirac equation? Despite its deep connection to the Klein-Gordon and Schrödinger equations, the Dirac system has a unique structure. Energy-based methods cannot be directly applied because the energy functional is not sign-definite. Furthermore, the Hamiltonian of the nonlinear Dirac equation is unbounded from above and below, reflecting the presence of both particles and antiparticles in the model [64]. Consequently, the energy does not provide a priori estimates. Concerning the existence of breather solutions in the 1D massless Gross–Neveu model, for large values of $N$ the existence of breather solutions under a symmetric scalar-scalar interaction was proved in [31]. In this context, breather solutions are interpreted as multi-fermion bound states oscillating in time at their rest frame. Recently, in [39], the authors described a method to find such solutions and studied the breather-breather scattering problem.
@@
It is interesting to discuss why Theorem 1.1 holds but it does not hold in the massive case. First of all, a simple dispersive analysis reveals that in the case $m=0$ linear waves escape to infinity with speed one, meaning that compact intervals of time remain free of these slowly decaying waves. This is not the case in the massive setting, which is more related to NLKG in that sense. We have found a particular new structure in the massless case that allows us to cover almost the whole light cone, with some loss coming from integrability estimates. This will not be the situation in the massive case.
@@
Let $(\psi_{1},\psi_{2})$ be any global solution in $C^{1}_{loc}(\mathbb{R};(L^{2}\times L^{2})(\mathbb{R};\mathbb{C})$ of the massless Dirac equation given in (1.8). Then, for $I(t)$ as in (1.9), there is a strong decay to zero on the $L^{2}$ norm. Therefore, no soliton nor breather solution exists for massless Dirac equation (1.8) inside the region $I(t)$, for any time t sufficiently large.
@@
The condition (1.12) is a “Cauchy-Riemann” type condition, and will ensure that solutions issued from odd data will continue having that property for late times, even if (1.11) seems do not respect that property in general. Therefore, the standard integrable nonlinearities and the odd-even or even-odd parities, naturally preserved by Dirac’s model, and the origin of the existence of non decaying solitary waves, will be discarded by this condition. As a consequence, one will have that $i\partial_{t}\psi_{1}-\partial_{x}\psi_{2}$ and $i\partial_{t}\psi_{2}+\partial_{x}\psi_{1}$ will be odd for all times. Additionally, under better smoothness conditions (1.12) imply the natural condition
@@
giving the reason on why we adopted the name “harmonic” (or Cauchy-Riemann related) property. Finally, condition (1.13) is necessary to get “conservation” of a suitable Lyapunov functional, meaning that more general nonlinearities may not be able to retain boundedness in time of solutions. For examples of nonlinearities satisfying these conditions, see Section 1.7. For this model, we have the following decay property:
@@
The proof of Theorem 1.3 involves the introduction of new virial identities adapted to the Dirac setting, without requiring to square the equation to get a NLKG model. Indeed, performing this classical trick may be more subtle than usual, see Appendix B for details. Indeed, it is proved in Lemma B.1 that the classical process of “squaring the equation” may be not completely rigorous unless one assumes more conditions on the data and the corresponding solutions. In our case, estimates are obtained in the purely Dirac form (1.11), recovering the dynamics of $\psi_{1}$ and $\psi_{2}$ at the same time. We believe that virial estimates of this type are new and introduce an interesting direction where one can further investigate in the future, in particular in the case of manifolds around solitary wave solutions.
Now we consider the 3D case. For the 3D magnetic linear Dirac equation, a potential-perturbed version of the linear Dirac equation, but not considered in this work, the authors in [15] derived a virial identity and subsequently used it to obtain smoothing and endpoint Strichartz estimates. Additionally, they proved a Hardy-type inequality for the perturbed Dirac operator. Due to the lack of a definite sign in the Dirac operator, the authors addressed this challenge by considering the squared Dirac equation. Later, in [17], the author extended these results to dimensions higher than three. The aforementioned works relied on a virial identity, which leveraged the hidden Klein-Gordon structure. Furthermore, the Hamiltonian ${\mathcal{H}}_{m}$, derived from the Dirac operator ${\mathcal{D}}_{m}$, is unitarily equivalent to the square root of the Klein-Gordon Hamiltonian $\gamma^{0}\sqrt{-\Delta+m^{2}}$. In other words, it satisfies ${\mathcal{H}}_{m}^{2}=(-\Delta+m^{2})I_{N}$ (see [64, 19, 17, 15], and the references therein.) Finally, see [21] for results on long time behavior of solutions in the case of linear and nonlinear Dirac models on manifolds.
@@
where the spaces $\mathcal{H}_{m_{j},k_{j}}$ are called partial wave subspaces, and each partial wave subspaces is of dimension two. The spaces $\mathcal{H}_{m_{j},k_{j}}$ is given by
@@
(See [64, Section 4.6].) In order to simplify the discussion, here we are just considered a subspace of dimension 1. This choose avoids the existence of bi-frequency solitary waves. In particular, let us consider the case of “spherical coordinates” $(r,\theta,\phi)\in(0,\infty)\times(0,\pi)\times(0,2\pi)$; and let us consider a solution of the Dirac equation of the form (Soler [62])
@@
Notice that with this choice we are working on a subspace of the partial wave space. Then, after some classical computations one gets
@@
Therefore, the 3D Dirac equation (1.3) as a $4\times 4$ complex system with $V=-g(u,v)$ associated to this partial wave subspace-type solution (1.16) is given by
@@
where for each $j=1,2$ we decompose $W_{j}=W_{j1}+iW_{j2}$, $W_{j1},W_{j2}$ real-valued. Assume additionally that
  • There exists ${\bf W}:\mathbb{C}^{4}\to\mathbb{R}$, ${\bf W}={\bf W}(\phi_{1},\overline{\phi}_{1},\phi_{2},\overline{\phi}_{2})$, such that $|{\bf W}|\leq C|\phi|^{p}$ with $p\geq 4$, $|\phi|=\sqrt{|\phi_{1}|^{2}+|\phi_{2}|^{2}}$, and
    $W=(W_{1},W_{2})=({\partial_{\overline{\phi}_{1}}{\bf W}},{\partial_{\overline{\phi}_{2}}{\bf W}}).$ (1.20)
Remark 1.3 (On Soler nonlinearities).
We remark that condition (1.20) precludes classical models with solitary waves such as the Soler nonlinearities. Indeed, in the Soler case
where $g$ is polynomial and $g(0)=0$, one has that there is no ${\bf W}$ satisfying (1.20) because of a wrong a sign condition in (1.18), meaning that in the Soler case one should have $(W_{1},W_{2})=({\partial_{\overline{\phi}_{1}}{\bf W}},-{\partial_{\overline{\phi}_{2}}{\bf W}})$.
@@
By a classical Sobolev embedding, one also has decay of local Sobolev subcritical $L^{p}$ norms. Notice that no condition on the parity of the data is required, and (1.12) is also not required in the 3D case. This is coincident with the fact that it is expected that breathers are difficult to obtain in dimensions bigger than one. Mathematically speaking, decay motivated ”virial identities” in the radial 3D case present better spectral properties than in the 1D case [49], making more quantitative the classical argument ”the bigger the dimension, the better the linear decay”. This is reflected in the fact that one can cook up suitable virial weights under which local decay is achieved, and this procedure fails in 1D by classical counterexamples (breathers). In this paper, after a suitable decomposition of $\phi$ in components $(\phi_{11},\phi_{12},\phi_{21},\phi_{22})$, we will consider 3D radial virial functionals of the form
@@
with $\varphi=\frac{r^{3/2}}{1+r}$, plus other two related virials $\mathcal{K}_{2}$ and $\widetilde{\mathcal{K}_{2}}$. The combination of these four functionals and the precise choice of the weight $\varphi$ with a strong decay at $r=0$ but an additional decay at infinity will help us to describe the long time local in space dynamics of small globally bounded 3D Dirac massless and massive waves, allowing us to prove (1.22).

### v4

@@ Abstract.
We study the long-time behavior of small and large solutions to a broad class of nonlinear Dirac-type equations. Our results are classified in 1D massless and massive cases, 3D general and $n$-dimensional in generality. In the 1D massless case, we prove that any globally defined solution converges to zero as time tends to infinity, within a spatial region expanding at a rate proportional to $t\log^{-2}t$. This result holds without assumptions on the smallness of initial data or specific power of nonlinearity, ruling out the existence of standing breather-like or solitary wave structures in this regime. In the 1D massive case, solitary waves are known to exist. Introducing new virial identities adapted to Dirac’s distinctive algebra, we prove that there are “holomorphic” odd nonlinearities under which globally defined small odd solutions decay to zero on spatial compact sets as time tends to infinity. This result is extended to the 3D case under boundedness of the $H^{1}$ norm but without requiring the parity condition on the data, giving decay proofs for an important class of nonlinear Dirac models, and opening the door to the future use of virial identities to prove asymptotic stability of well-chosen Dirac solitary waves.
@@ 1.1. Setting
The (linear) Dirac equation was introduced as a relativistic version of the Schrödinger equation [32], being one of the most relevant models in relativistic quantum mechanics. It describes the self-interaction of Dirac fermions [32, 64]. Several authors have introduced and justified a nonlinear version of this model, see e.g., Soler and Thirring models [62, 65, 1]. Generally, the nonlinear Dirac model is
@@
Here $\psi:\mathbb{R}\times\mathbb{R}^{n}\to\mathbb{C}^{N}$ is a spinor-valued wave function, $m\in\mathbb{R}$ is the mass, and $N=2^{\lfloor(n+1)/2\rfloor}$, where the summation convention is $\gamma^{\mu}\partial_{\mu}=\sum_{\mu=0}^{n}\gamma^{\mu}\partial_{\mu}$, and $\partial_{0}\equiv\partial_{t}$. The Dirac matrices $\gamma^{\mu}$ are chosen such that they satisfy the anti-commutativity property
@@
and where $\eta^{\mu\nu}$ is related to the Minkowski metric $\eta=\mbox{diag}(1,-1,\dots,-1)$. Some members of the Dirac family equation are invariant under Lorentz boosts, e.g., the Soler and Thirring models.
@@ 1.2. Well-posedness
The well-posedness problem of the nonlinear Dirac equation is challenging due to the lack of a defined sign on the Hamiltonian. Most works address the massive and the massless cases separately. The literature is extensive across both 1D and higher dimensions; here, we highlight the most relevant works.
In the 1D case, Candy [18] used null coordinates to prove global existence in $H^{s}$ for all $s\geq 0$ in the case of the $L^{2}$-critical and integrable massive Thirring model. Pelinovsky [58] showed the existence of small-$H^{1}$-norm solutions for more general nonlinear Dirac equations with cubic and higher-order nonlinear terms, and also established scattering for global solutions with small initial data in $H^{1}\cap L^{1}$. In [46] the authors proved local well-posedness in $H^{s}$ with s>1/2s>-1/2 and ill-posedness in $H^{-1/2}$ for the quadratic nonlinear Dirac equation. See [18, 58, 46, 54, 59, 7, 10, 60] and references therein for further details.
@@ 1.3. Main results
In this paper, our main goal is to study the long-time asymptotic behavior of globally defined nonlinear Dirac solutions, under minimal assumptions, and including integrable and nonintegrable models. Highly relevant for us will be the 1D and 3D cases. We first consider the 1D Dirac equation in laboratory coordinates:
@@
$U_{\omega}(x)=\frac{\gamma}{\sqrt{1+\omega}\cosh(\gamma x)+i\sqrt{1-\omega}\sinh(\gamma x)},\quad\gamma=\sqrt{1-\omega^{2}}.$
These (standing) solitary waves are a clear obstruction to decay, having interesting symmetries. Indeed, one can see that among the symmetries preserved by (1.6) one can find the (even + $i$ odd) and (odd + $i$ even) ones, which are precisely present in the solitary waves (1.7). Moreover, notice that the classical (odd + $i$ odd) symmetry, natural in scalar field models with odd nonlinearities, is not preserved this time. This makes the exercise of finding decay in the energy space in the 1D case a highly complicated task. Indeed, we shall concentrate efforts in another manifold of initial data not producing solutions such as (1.7). Within this region, we shall obtain energy local decay.
@@
$H(u,v)=$ $~{}\Im\int(u_{x}\overline{u}-v_{x}\overline{v})dx+\Re\int(u\overline{v}-W(u,v))dx,$
$P(u,v)=$ $~{}\Im\int(u\overline{u}_{x}+v\overline{v}_{x})dx,\qquad Q(u,v)=\int(|u|^{2}+|v|^{2})dx,$
@@
It is well-known that the cubic Nonlinear Schrödinger (NLS) equation admits breather solutions. These solutions are characterized by being nontrivially spatially localized and periodic in time, distinguishing them from standing waves. Breather solutions possess a more intricate analytical structure; for example, they may involve complex interactions between different modes, which result in oscillatory behavior that is both spatially and temporally localized. One of the well-known breathers for the focusing cubic NLS is the Satsuma-Yajima breather solution (see [4, 2] and the reference therein). For the Klein-Gordon equation (and for a class of 1D nonlinear wave equations), in [44], the authors proved the nonexistence of small odd breather solutions (see also [61]).
Similarly, Martínez [48] showed that, for a wide range of nonlinearities, any global odd solution to the 1D NLS equation decays to zero in compact regions of space as time tends to infinity, assuming only initial data in the energy space. Previous work addressed the decay/scattering problem under the condition that the power $p$ satisfies $3<p<5$, but this result improved the range, including $1<p\leq 3$. More recently, [2] characterized the nonexistence of breather solutions for the $n$-dimensional NLS using the energy and power of the corresponding nonlinearity.
Given the similarities between the Dirac equation and both the Klein-Gordon and the classical Nonlinear Schrödinger equations, a natural question arises: Do breather solutions exist for the Dirac equation? Despite its deep connection to the Klein-Gordon and Schrödinger equations, the Dirac system has a unique structure. Energy-based methods cannot be directly applied because the energy functional is not sign-definite. Furthermore, the Hamiltonian of the nonlinear Dirac equation is unbounded from above and below, reflecting the presence of both particles and antiparticles in the model [64]. Consequently, the energy does not provide a priori estimates. Concerning the existence of breather solutions in the 1D massless Gross–Neveu model, for large values of $N$, the existence of breather solutions under a symmetric scalar-scalar interaction was proved in [31]. In this context, breather solutions are interpreted as multi-fermion bound states oscillating in time at their rest frame. Recently, in [39], the authors described a method to find such solutions and studied the breather-breather scattering problem.
@@
It is interesting to discuss why Theorem 1.1 holds, but it does not hold in the massive case. First of all, a simple dispersive analysis reveals that in the case $m=0$ linear waves escape to infinity with speed one, meaning that compact intervals of time remain free of these slowly decaying waves. This is not the case in the massive setting, which is more related to NLKG in that sense. We have found a particular new structure in the massless case that allows us to cover almost the whole light cone, with some loss coming from integrability estimates. This will not be the situation in the massive case.
@@
Let $(\psi_{1},\psi_{2})$ be any global solution in $C^{1}_{loc}(\mathbb{R};(L^{2}\times L^{2})(\mathbb{R};\mathbb{C})$ of the massless Dirac equation given in (1.8). Then, for $I(t)$ as in (1.9), there is a strong decay to zero on the $L^{2}$ norm. Therefore, no soliton nor breather solution exists for the massless Dirac equation (1.8) inside the region $I(t)$, for any time t sufficiently large.
@@
The condition (1.12) is a “Cauchy-Riemann” type condition, and will ensure that solutions issued from odd data will continue having that property for late times, even if (1.11) seems do not respect that property in general. Therefore, the standard integrable nonlinearities and the odd-even or even-odd parities, naturally preserved by Dirac’s model, and the origin of the existence of non-decaying solitary waves, will be discarded by this condition. As a consequence, one will have that $i\partial_{t}\psi_{1}-\partial_{x}\psi_{2}$ and $i\partial_{t}\psi_{2}+\partial_{x}\psi_{1}$ will be odd for all times. Additionally, under better smoothness conditions (1.12) implies the natural condition
@@
giving the reason why we adopted the name “harmonic” (or Cauchy-Riemann related) property. Finally, condition (1.13) is necessary to get “conservation” of a suitable Lyapunov functional, meaning that more general nonlinearities may not be able to retain boundedness in time of solutions. For examples of nonlinearities satisfying these conditions, see Section 1.7. For this model, we have the following decay property:
@@
The proof of Theorem 1.3 involves the introduction of new virial identities adapted to the Dirac setting, without requiring to square of the equation to get a NLKG model. Indeed, performing this classical trick may be more subtle than usual, see Appendix B for details. Indeed, it is proved in Lemma B.1 that the classical process of “squaring the equation” may not be completely rigorous unless one assumes more conditions on the data and the corresponding solutions. In our case, estimates are obtained in the purely Dirac form (1.11), recovering the dynamics of $\psi_{1}$ and $\psi_{2}$ at the same time. We believe that virial estimates of this type are new and introduce an interesting direction where one can further investigate in the future, in particular in the case of manifolds around solitary wave solutions.
Now we consider the 3D case. For the 3D magnetic linear Dirac equation, a potential-perturbed version of the linear Dirac equation, but not considered in this work, the authors in [15] derived a virial identity and subsequently used it to obtain smoothing and endpoint Strichartz estimates. Additionally, they proved a Hardy-type inequality for the perturbed Dirac operator. Due to the lack of a definite sign in the Dirac operator, the authors addressed this challenge by considering the squared Dirac equation. Later, in [17], the author extended these results to dimensions higher than three. The aforementioned works relied on a virial identity, which leveraged the hidden Klein-Gordon structure. Furthermore, the Hamiltonian ${\mathcal{H}}_{m}$, derived from the Dirac operator ${\mathcal{D}}_{m}$, is unitarily equivalent to the square root of the Klein-Gordon Hamiltonian $\gamma^{0}\sqrt{-\Delta+m^{2}}$. In other words, it satisfies ${\mathcal{H}}_{m}^{2}=(-\Delta+m^{2})I_{N}$ (see [64, 19, 17, 15], and the references therein.) Finally, see [21] for results on long-time behavior of solutions in the case of linear and nonlinear Dirac models on manifolds.
@@
where the spaces $\mathcal{H}_{m_{j},k_{j}}$ are called partial wave subspaces, and each partial wave subspace is of dimension two. The spaces $\mathcal{H}_{m_{j},k_{j}}$ is given by
@@
(See [64, Section 4.6].) In order to simplify the discussion, here we are just considering a subspace of dimension 1. This choice avoids the existence of bi-frequency solitary waves. In particular, let us consider the case of “spherical coordinates” $(r,\theta,\phi)\in(0,\infty)\times(0,\pi)\times(0,2\pi)$; and let us consider a solution of the Dirac equation of the form (Soler [62])
@@
Notice that with this choice, we are working on a subspace of the partial wave space. Then, after some classical computations, one gets
@@
Therefore, the 3D Dirac equation (1.3) as a $4\times 4$ complex system with V=g(u,v)V=-g(u,v) associated to this partial wave subspace-type solution (1.16) is given by
@@
where for each $j=1,2$ we decompose $W_{j}=W_{j1}+iW_{j2}$, $W_{j1},W_{j2}$ real-valued.
Additionally, consider that the nonlinearity is of pure power type. This assumption allows us to capture a wide range of nonlinearities with physical relevance. Let $(W_{1},W_{2})$ be a nonlinearity such that
  • There exists $C>0$ such that for $(W_{1},W_{2})=(W_{1},W_{2})(\phi_{1},\overline{\phi}_{1},\phi_{2},\overline{\phi}_{2})$
    $|W_{1}|+|W_{2}|\leq C|\phi|^{p}\mbox{ with }p\geq 3.$ (1.20)
@@
Remark 1.3 (On Soler’s nonlinearities).
Recall that condition (1.20) includes classical models with solitary waves, such as the Soler’s nonlinearities. Indeed, in the Soler’s case one has
where $g$ is a polynomial with $g(0)=0$. It follows that $(W_{1},W_{2})$ satisfies (1.20). In other words, for Soler’s nonlinearity, the above Theorem holds, implying that there are no arbitrarily small and localized solitary waves or breather solutions.
By a classical Sobolev embedding, one also has decay of local Sobolev subcritical $L^{p}$ norms. Notice that no conditions on the parity of the data are required, and (1.12) is also not required in the 3D case. This is consistent with the fact that it is expected that breathers are difficult to obtain in dimensions larger than one. Mathematically speaking, decay motivated ”virial identities” in the radial 3D case present better spectral properties than in the 1D case [49], making more quantitative the classical argument ”the bigger the dimension, the better the linear decay”. This is reflected in the fact that one can cook up suitable virial weights under which local decay is achieved, and this procedure fails in 1D by classical counterexamples (breathers). In this paper, after a suitable decomposition of $\phi$ into components $(\phi_{11},\phi_{12},\phi_{21},\phi_{22})$, we will consider 3D radial virial functionals of the form
@@
with $\varphi=\frac{r^{3/2}}{1+r}$, plus other two related virials $\mathcal{K}_{2}$ and $\widetilde{\mathcal{K}_{2}}$. The combination of these four functionals and the precise choice of the weight $\varphi$ with a strong decay at $r=0$ but an additional decay at infinity will help us to describe the long-time local in space dynamics of small globally bounded 3D Dirac massless and massive waves, allowing us to prove (1.21).